# Let’s Go Shopping (LGS) – Web-Scale Image-Text Dataset for Visual Concept Understanding

### 09 Jan 2024 v1

@@ 1 Introduction
Finally, for text-to-image generation tasks, diffusion models [54, 20, 2] is currently the most popular family of methods. To illustrate the efficacy of LGS in this setting, we use Stable Diffusion (SD) [49] and fine-tune it in both general and fine-grained settings on subsets of the LGS dataset. We demonstrate promising qualitative and quantitative results on adapting existing text-to-image models using LGS for e-commerce-related generations. Furthermore, with the help of its distinct image style and descriptive captions, LGS can help the SD model generate e-commerce-styled images.
@@ 4.1.2 Non-classification Visual Feature Extractors Can Generalize
To quantitatively demonstrate the generalizability of the vision feature extractors, we evaluate the reconstruction performance of the MAE models trained on LGS and ImageNet on COCO. The qualities of the raw reconstructions obtained by the models are presented in Table 5. While LGS is more domain-specific compared with ImageNet and COCO (both of which cover a wide range of domains), the MAE trained on LGS is able to generate COCO images with higher qualities compared with the ImageNet model. Furthermore, we use Table 6 to show that upon the visual embeddings learned jointly on ImageNet and LGS, a linear classifier with a satisfactory performance can be learned on both ImageNet and LGS. The above results verify that the feature extractors can generalize between LGS and general-domain datasets, despite the separation of the intermediate visual embeddings (which are visualized in Section A.2).
@@ 4.1.3 LGS Supplements ImageNet as a Pre-training Dataset
In Section A.3 in the supplementary materials, we use GradCam [14, 52] to visualize the representations learned by the classification models, demonstrating that the LGS models look for much more localized patterns that are relevant to e-commerce classification.
@@ 4.3 Text-to-Image Generation
For the general setting, we add a domain identifier to all training prompts associated with LGS images and guide the SD model to adapt to the e-commerce image style when this identifier is provided. The choice of the domain identifier is crucial, as the paper [50] shows that a domain identifier with a strong prior should be avoided. For example, the word retail has a strong prior, and the pre-trained “Vanilla” SD model confidently associates it with (physical) retail stores. This behavior is undesirable for the goal of e-commerce style transfer. By analyzing the effects of various domain identifiers on the generations of the pre-trained SD model, we determine that the word “e-commerce” gives a weak prior and is a suitable identifier. We then construct the ground-truth training prompts for the LGS images in the format of a photo of <brand> <end_leaf> <title>, e-commerce, where the <end_leaf> refers to the end leaf of the taxonomy tree introduced in Section 3.4. The “Vanilla” SD checkpoint is fine-tuned on one million LGS image-prompt pairs for 100k steps with a batch size of 24. Table 9 displays the quantitative results on an unseen validation set (5K image-prompt pairs) from LGS and a subset of the DeepFashion InShop dataset. The fine-tuning process enhances the performance of SD on LGS as expected. While the FID scores on DeepFashion are lower, the generations of the LGS-117 fine-tuned model are aesthetically more appealing. At this instant, there are no quantitative metrics that directly measure aesthetic superiority. Thus, we present Figure 6 and the additional examples in Section A.6 in the supplementary materials (Figures Supp-9 and Supp-8) to demonstrate the aesthetic improvement qualitatively. The lower FID scores may indicate a distribution shift between LGS and DeepFashion images.
@@ 6 Acknowledgment
Appendix A Additional Analyses
@@ A.1 LGS End Leaf Histogram
The instance counts of the LGS end leaves are displayed in Figure Supp-1. The top 80 most popular end leaves encompass 83.28% of the total instances, with the most popular Tops and T-shirts containing 16.23% of the instances.
A.2 How Features learned on ImageNet and LGS Differ
@@
A.3 LGS Classification Models Look for Localized Patterns
@@
A.4 Linear Probing Details
In this section of the appendix, we discuss the implementation details for the linear probing experiments in Section 4.1.3. In the existing literature, when ResNets (designed for $224\times 224$ inputs) are adopted for tasks that use smaller input sizes, the first $7\times 7$ convolution layer is often replaced with a $3\times 3$ layer. We adopt this replacement for CIFAR and Fashion MNIST. During linear probing, we thus allow this modified, randomly reinitialized first layer to be optimized along with the output layer.
@@ A.4 Linear Probing Details
A.5 $n$-gram and POS Analysis of LGS Captions
Figure Supp-5: Top 20 most common words per POS for LGS.
Figure Supp-6: Top 20 most common words per POS for COCO.
Number of $n$-grams with occurrence $\geq 10$   Five most frequent $n$-grams ($n=1,2,3$)
          LGS COCO   LGS COCO
uni-grams   364,802 17,009   and, the, a, to, with a, of, on, the, i
bi-grams   4,054,418 184,882   with a, in the, of the, is a, for a on a, in a, a man, of a, with a
tri-grams   8,900,084 462,653   true to size, made to order, this is a, a group of, group of people
this item is, machine wash cold in front of, next to a , on top of
A.6 Determining the Prompts for Text-to-Image Generation
@@
Quantitatively, Prompts 3 and 4 perform significantly better on LGS, perform comparably on DeepFashion, and generalize well (Table Supp-2). Prompt 3 had better FID scores using the Vanilla model and performed slightly better on LGS. Qualitatively, however, Prompt 4 generations are consistently better and more faithful to the metadata (Figure Supp-7). Therefore, we choose Prompt 4 for our experiments. This also reaffirms that these metrics are not strong indicators of aesthetic quality in this particular case, and should only be used as a loose relative measure. Figures Supp-9 and Supp-8 show additional examples from the two datasets generated with Prompt 4.

### 05 Mar 2024 v2

@@ 1 Introduction
Finally, for text-to-image generation tasks, diffusion models [54, 20, 2] form the currently most popular family of methods. To illustrate the efficacy of LGS in this setting, we use Stable Diffusion (SD) [49] and fine-tune it in both general and fine-grained settings on subsets of the LGS dataset. We demonstrate promising qualitative and quantitative results on adapting existing text-to-image models using LGS for e-commerce-related generations. Furthermore, with the help of its distinct image style and descriptive captions, LGS can help the SD model generate e-commerce-styled images.
@@ 4.1.2 Non-classification Visual Feature Extractors Can Generalize
To quantitatively demonstrate the generalizability of the vision feature extractors, we evaluate the reconstruction performance of the MAE models trained on LGS and ImageNet on COCO. The qualities of the raw reconstructions obtained by the models are presented in Table 5. While LGS is more domain-specific compared with ImageNet and COCO (both of which cover a wide range of domains), the MAE trained on LGS is able to generate COCO images with higher qualities compared with the ImageNet model. Furthermore, we use Table 6 to show that upon the visual embeddings learned jointly on ImageNet and LGS, a linear classifier with a satisfactory performance can be learned on both ImageNet and LGS. The above results verify that the feature extractors can generalize between LGS and general-domain datasets, despite the separation of the intermediate visual embeddings (which are visualized in Section B.1).
@@ 4.1.3 LGS Supplements ImageNet as a Pre-training Dataset
In Section B.2 in the supplementary materials, we use GradCam [14, 52] to visualize the representations learned by the classification models, demonstrating that the LGS models look for much more localized patterns that are relevant to e-commerce classification.
@@ 4.3 Text-to-Image Generation
For the general setting, we add a domain identifier to all training prompts associated with LGS images and guide the SD model to adapt to the e-commerce image style when this identifier is provided. The choice of the domain identifier is crucial, as the paper [50] shows that a domain identifier with a strong prior should be avoided. For example, the word retail has a strong prior, and the pre-trained “Vanilla” SD model confidently associates it with (physical) retail stores. This behavior is undesirable for the goal of e-commerce style transfer. By analyzing the effects of various domain identifiers on the generations of the pre-trained SD model, we determine that the word “e-commerce” gives a weak prior and is a suitable identifier. We then construct the ground-truth training prompts for the LGS images in the format of a photo of <brand> <end_leaf> <title>, e-commerce, where the <end_leaf> refers to the end leaf of the taxonomy tree introduced in Section 3.4. The “Vanilla” SD checkpoint is fine-tuned on one million LGS image-prompt pairs for 100k steps with a batch size of 24. Table 9 displays the quantitative results on an unseen validation set (5K image-prompt pairs) from LGS and a subset of the DeepFashion InShop dataset. The fine-tuning process enhances the performance of SD on LGS as expected. While the FID scores on DeepFashion are lower, the generations of the LGS-117 fine-tuned model are aesthetically more appealing. At this instant, there are no quantitative metrics that directly measure aesthetic superiority. Thus, we present Figure 6 and the additional examples in Section C.1 in the supplementary materials (Figures Supp-9 and Supp-8) to demonstrate the aesthetic improvement qualitatively. The lower FID scores may indicate a distribution shift between LGS and DeepFashion images.
@@ 6 Acknowledgment
Appendix A Additional Analyses on LGS’s Data Distribution
@@ A.1 LGS End Leaf Histogram
The instance counts of the LGS end leaves are displayed in Figure Supp-1. The top 80 most popular end leaves encompass 83.28% of the total instances, with the most popular Tops and T-shirts containing 16.23% of the total instances.
A.2 $n$-gram and POS Analysis of LGS Captions
Figure Supp-5: Top 20 most common words per POS for LGS.
Figure Supp-6: Top 20 most common words per POS for COCO.
Number of $n$-grams with occurrence $\geq 10$
Five most frequent $n$-grams ($n=1,2,3$)
          LGS COCO   LGS COCO
uni-grams   364,802 17,009   and, the, a, to, with a, of, on, the, i
bi-grams   4,054,418 184,882
with a, in the, of the, is a, for a
on a, in a, a man, of a, with a
tri-grams   8,900,084 462,653
true to size, made to order, this is a,
a group of, group of people
this item is, machine wash cold
in front of, next to a, on top of
Appendix B Additional Analyses on LGS-trained Classifiers
B.1 How Features learned on ImageNet and LGS Differ
@@
B.2 LGS Classification Models Look for Localized Patterns
@@
B.3 Linear Probing Details
In this section, we discuss the implementation details for the linear probing experiments in Section 4.1.3. In the existing literature, when ResNets (designed for $224\times 224$ inputs) are adopted for tasks that use smaller input sizes, the first $7\times 7$ convolution layer is often replaced with a $3\times 3$ layer. We adopt this replacement for CIFAR and Fashion MNIST. During linear probing, we thus allow this modified, randomly reinitialized first layer to be optimized along with the output layer.
@@ B.3 Linear Probing Details
Appendix C Additional Text-to-Image Generation Discussions
C.1 Determining the Prompts for Text-to-Image Generation
@@
Quantitatively, Prompts 3 and 4 perform significantly better on LGS, perform comparably on DeepFashion, and generalize well (Table Supp-2). Prompt 3 achieves better FID scores using the Vanilla model and performs slightly better on LGS. Qualitatively, however, Prompt 4 generations are consistently better and more faithful to the metadata (Figure Supp-7). Therefore, we select Prompt 4 for our experiments. This also reaffirms that these objective metrics are not strong indicators of the subjective aesthetic quality in this particular case, and should only be used as a loose relative measure. Figures Supp-9 and Supp-8 show additional examples from the two datasets generated with Prompt 4.